INFINITE DIMENSIONAL FAMILIES OF LOCALLY NONSOLVABLE PARTIAL DIFFERENTIAL OPERATORS

MICHAEL CHRIST GEORGI KARADZHOV DETLEF MÜLLER

Abstract. Local solvability is analyzed for natural families of partial differential operators having double characteristics. In some families the set of all operators that are not locally solvable is shown to have both infinite dimension and infinite codimension.